40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5383     Entry time: Sat Sep 10 20:30:01 2011
Author: rana 
Type: Update 
Category: IOO 
Subject: MC trans re-aligned / MC2 shifted mysteriously / MC2 re-aligned 

MCdecenter.png

I re-aligned the beam onto the MC TRANS QPD since Kiwamu had centered the spots on the mirrors. However, I then inspected the MC2F camera. After coming back into the control room I noticed that the MC transmission had gone down by 50% and that the MC2 OSEMs showed a large step. My guess is that somehow the opening and closing of the can shifted the suspension. So I adjusted the MC2 alignment biases to recover the transmitted power (its now ~50000 instead of the ~33000 from Friday).

ELOG V3.1.3-